# A Digital twin for Diesel Engines: Operator-infused PINNs with Transfer Learning for Engine Health Monitoring

### v1

@@ 2.1 Main concept
    Empirical relationships: First, we consider various empirical relationships between variables in the engine model. In our previous study [20], we demonstrated how these empirical formulae can be approximated using deep neural networks (DNN). The empirical formulae are engine-specific, and DNNs can provide a generalized representation when the exact relationships are unknown by training the DNN with data from the specific engine. In addition, to account for the systematic (e.g. manufacturing) variations of the engine, we introduce stochasticity into the empirical relationship by introducing dropout in the DNN.
@@
    Independent States Approximation: Second, we note here that the states describing the EGR and VGT dynamics are independent of the rest of the states, with no unknown parameters in these dynamics. We approximated these states using two DeepONets, one for the two EGR states ($\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$) and one for the VGT state ($\tilde{u}_{vgt}$) as discussed in section 2.3. These DeepONets are pre-trained using simulation data and used as surrogates to predict the actuator states.
@@
    Hybrid method: Third, we propose a hybrid method combining the pre-trained DNN, the pre-trained DeepONet and PINN model, as detailed in section 2.4.
@@
    Transfer Learning Strategies: Lastly, to reduce the computation cost (in terms of the number of epochs and trainable parameters), we consider two transfer learning strategies:
@@
      a hybrid PINN multi-stage transfer learning method, discussed in section 2.5, and
@@
    We also use self-adaptive weights [26] for better and faster convergence in all examples of the proposed methods.
Furthermore, as noted above, we have pre-trained networks for empirical formulae (DNNs) and for EGR and VGT states (DeepONets). These networks are trained using simulated data (we label this as laboratory data for nomenclature simplicity), which corresponds to the design value of the unknown parameters. However, in the field, the values of the unknown parameters may change due to natural systematic changes over time. Thus, we test our proposed hybrid PINN method on a set of parameters different from the design parameters used for generating laboratory data for training our empirical networks and DeepONets. This helps us validate whether our hybrid PINN model can accurately identify parameters independently of the predictions from the pre-trained networks, thus allowing us to analyze any potential bias originating from this hybrid approach. The different values of the unknown parameters considered for generating laboratory and field data are shown in Table 1.
Table 1: True value of estimated parameters considered for generating laboratory and field data conditions. The laboratory data is considered for training the DNNs for empirical formulae, DeepONets for EGR and VGT actuator dynamics, and training the body for the few-shot transfer learning. The field data represents data that is used for training the hybrid PINN framework for parameter estimation. For confidentiality of data, we have shown the true values for field data as 1, and the laboratory data are scaled with respect to the field data.
$A_{egramx}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$ Remarks
Laboratory data $1.333$ $1.1663$ $0.9179$ $1$ Considered for data generation to train the pre-trained networks and DeepONet
Field data $1$ $1$ $1$ $1$ Considered for data generation for field data (use in PINN inverse problem)
Mask and scaling consider in PINN
Mask function $\exp(.)$ $\exp(.)$ softplus(.) $\exp(.)$ This ensures the unknown parameters are always positive
Scaling ×$\times 10^{-4}$$\times 10^{-4}$$\times 1$$\times 10$ Scale to obtain the parameters in physical domain
@@ 2.2 Neural network surrogates for empirical formulae
The empirical equations are discussed in Appendix A
@@
where the subscript $i=1,2,\dots,6$ indicate different neural networks shown in Table 2, $n_{i}$ represents the number of labelled training data for $i$th neural network. $\bm{x}_{i}$ denotes the input corresponding to the $i$th network, $\hat{y}_{i}$ and $y_{i}$ represent the estimated output and the corresponding labelled values, respectively for the $i$th network. The laboratory data used for training each neural network have been discussed in our previous work [20] along with detailed calculations for each labelled data. We also consider a 3% Gaussian noise in the simulated data of the output variables during training to account for the field measurement conditions. These networks are trained using the Adam optimizer. After training, we integrate these DNNs into our hybrid PINN framework to replace the empirical models. These pre-trained neural networks are shown with the output $g(\hat{y})$ in Fig. 2 and Fig. 4 in sections 2.4 and 2.6.
@@ 2.3 DeepONets for EGR and VGT states
Based on the universal approximation theorem of operators [27], Lu et al. [22] proposed the Deep Operator Network (DeepONet), which maps an infinite-dimensional function ($u\in\bm{U}$) to another infinite dimensional output function ($v\in\bm{V}$) in the Banach space. The conventional DeepONet consists of two DNNs, known as branch and trunk networks. The trunk network takes the spatio-temporal information (e.g. time $t$, domain points ${x,y,z}$) as input and provides the bases function as output. The branch network takes the input function ($u\in\bm{U}$) at fixed points and approximates the coefficients for each basis function. Importantly, the DeepONet architecture does not impose restrictions on the type of networks used for trunk and branch networks. The ability of DeepONet to learn complex mapping between input and output functional spaces makes it suitable for various applications in scientific machine learning. Mathematically, DeepONet operates as an operator that maps $\mathcal{G}_{\bm{\theta}}:\bm{U}\rightarrow\bm{V}$, and the output is given as
@@
We note here that the engine is a continuous dynamical system, and this makes it challenging to identify proper operating sub-spaces for input-output mapping for the operator network learning. Hence, it is important to study the driving behaviour (engine running condition) over large duration and sample sizes. An appropriate window size (time duration) is required in which correct input-output function spaces can be assumed while preserving the dynamic behaviour of the system. Here, we consider a 60-sec window and approximate the independent state variables using Causality-DeepOnet [29], as proposed by Liu et al. [29]. Causality-DeepONet integrates causality and convolution of the input function in its formulation, which enhances accuracy in approximating oscillatory dynamic problems with a smaller training sample set. We find that this framework is particularly useful in capturing the rapidly changing dynamics of a diesel engine signal.
In this study, we use Causality-DeepONets to approximate the dynamics of the EGR actuator, i.e., $\tilde{u}_{egr1}$ (Eq. (6)) and $\tilde{u}_{egr2}$ (Eq. (7)) and the VGT actuator $\tilde{u}_{vgt}$ (Eq. (8)). Details of the two DeepONets are shown in Table 3. From hereon, we refer to Causality-DeepONet simply as DeepONet unless stated otherwise. These DeepONets approximate solutions to non-homogeneous initial value problems by learning the operator mapping between input function ($u$) and the initial condition ($\tilde{u}_{0}$) to the output ($\tilde{u}(t))$ using individual branch networks for the input function and initial condition, as described in Table 3. We note that the equations governing EGR and VGT actuator dynamics are first-order differential equations that can be computed efficiently using fast numerical methods. However, discrepancies between the mathematical model and experimental data were observed in [21]. By employing DeepONets trained on historical data, we obtain a better representation of real system dynamics as compared to a numerical method that relies on equations provided. Hence, our approach of using DeepONet surrogates for predicting actuator dynamics is likely to show better correlation when used on real system.
@@
DeepONet Variable Eq. No. of branch Branch inputs ($\bm{x}$)
$\mathcal{D}_{1}(t,\bm{x};\bm{\theta}_{DP1})$$\{\tilde{u}_{egr1}$, $\tilde{u}_{egr2}\}$ (A.16) and (A.17) 4 $\{u_{degr}$, $\tilde{u}_{egr1,0}\}$, $\{u_{degr}$, $\tilde{u}_{egr2,0}\}$
$\mathcal{D}_{2}(t,\bm{x};\bm{\theta}_{DP2})$$\{\tilde{u}_{vgt}\}$ (A.25) 2 $u_{dvgt}$, $\tilde{u}_{vgt,0}$,
Thus, we consider the operator mapped from the input function ($u$) and the initial condition ($\tilde{u}_{0}$) to the output ($\tilde{u}(t))$. We consider multiple branch nets for the input function and initial value as illustrated in Table 3. We also like to note that the equations describing the EGR and VGT states are first-order differential equations. Thus, it is possible to implement a fast numerical algorithm for their solutions. However, it was observed in [21] that the mathematical model and experimental data show slight disparity when an ensembled mean is considered. By considering DeepONet trained with historical data, we can obtain an average model of the output, which is a better representation than the numerical model.
$\mathcal{L}_{i}(\bm{\theta}_{DP,i})=$ $\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\hat{y}_{i}^{(j)}\right]^{2},\;\;\;\;\;i=1,2$ (14a)
$=$ $\dfrac{1}{n_{i}}\sum_{j=1}^{n_{i}}\left[y_{i}^{(j)}-\mathcal{D}_{i}(t,\bm{x}_{i};\bm{\theta}_{DP,i})^{(j)}\right]^{2},\;\;\;\bm{x}\rightarrow\text{Branch inputs}$ (14b)
We also note that the DeepONet for the EGR states has two output quantities. Consequently, the loss function for this DeepONet is the weighted summation of the loss function for individual output quantities. The network architecture, loss function and input-output are discussed in more detail in section 3, along with specific examples. These networks are trained using Adam optimizer. As discussed earlier, we utilize simulation data to train the DeepONets and subsequently replace the EGR and VGT dynamics with these pre-trained networks within the PINN model. Note that for DeepONet training here, we do not incorporate dropouts since the outputs of these DeepONets are used as inputs to the pre-trained empirical networks that include dropouts to inherently addresses the variability, as discussed in section 2.2. In the following section, we present details about our hybrid PINN framework to tackle the parameter estimation problem.
@@ 2.4 A hybrid model using DeepONet and PINN for the engine problem
In the previous section (2.3), we discussed how the three states corresponding to EGR and VGT can be approximated using DeepONet surrogates. In section 2.2, we introduced how we use DNNs to estimate empirical relationships in the engine model. In this section, we will discuss in detail how we construct a hybrid model combining DeepONet, empirical DNNs, and PINN to solve inverse problems. We propose a hybrid PINN model combining DeepONet and DNN to predict the gas flow dynamics and estimate unknown parameters.
@@
where $\bm{\theta}$ denotes the set of parameters of the neural network, i.e. the weights, $\bm{W}$ and the biases, $\bm{b}$; $\bm{\mathcal{\varLambda}}$ are the unknown scalar parameters to be estimated. The parameters $\lambda_{1}$ and $\lambda_{2}$ are the weights which control the convergence of each loss component. These $\lambda$s values may be constant or self-adaptive [26] depending on the problem setup and solution method. The data loss may be considered as the mean squared error (MSE) loss of the known labelled data, given by:
@@
where $N$ is the number of collocation points chosen for solving the equation and $r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})$ represents the residual of the corresponding differential equation. The residual $r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})$ should ideally be zero throughout the entire domain. The objective, therefore, is to obtain the optimal parameters ($\bm{\theta}$) of the neural network as well as the unknown parameters ($\bm{\mathcal{\varLambda}}$) in the system. We minimize the loss function Eq. (15) to obtain the optimal values of these parameters. Furthermore, we also incorporate self-adaptive weights [26] in the loss. We minimize the loss function with respect to the parameters ($\bm{\theta}$ and $\bm{\varLambda}$) and maximize with respect to the self-adaptive weights ($\lambda$). Thus, the optimization process may be written as,
@@
$\bm{\theta},\bm{\varLambda}$ $=\bm{\theta},\bm{\varLambda}-lr_{\theta,\varLambda}\nabla_{\theta,\varLambda}\mathcal{L}(\bm{\theta},\varLambda)$ (19a)
$\lambda$ $=\lambda+lr_{\lambda}\nabla_{\lambda}\mathcal{L}(\bm{\theta},\varLambda),$ (19b)
where $lr_{\theta}$ denotes the network optimiser learning rate for $\bm{\theta}$, $lr_{\varLambda}$ represents the learning rate for the optimizer used to identify unknown scalar parameters $\bm{\varLambda}$, and $lr_{\lambda}$ refers to the learning rate for the optimizer used for updating self-adaptive weights. Incorporating distinct learning rates and optimizers allows for fine tuning of each component in this hybrid framework for efficient convergence.
@@
To integrate the complete hybrid model, we combine the DNN models used for approximating the empirical formulae as discussed in section 2.2 and the two DeepONet models for approximating actuator states $\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$ with the PINN DNN model, as shown in figure 2. The parameters of these DeepONets and DNNs (for the empirical formulae) are kept fixed when we train the hybrid PINN model. Furthermore, in Eq. (15), we demonstrated how PINNs can be used for solving a single differential equation. For systems involving multiple equations, one can either use one DNN with multiple outputs or multiple DNNs each with one output. The loss function in Eq. (15) can be extended to include the loss function for all the equations and data by summing up individual loss functions for each equation and data. This hybrid approach leverages pre-trained networks and a PINN network that can be trained on live data for efficient and accurate estimation of engine parameters and state variables.
@@
Fig. 2: Schematic of a hybrid model of DeepONet and physics-informed neural networks (PINNs) for inverse problems. In the left part of the figure (enclosed in red dashed line), a PINN approximates the solution $y$ to a differential equation with time $t$ as input. The top left section (enclosed within black hatched lines) represents another DNN which takes as input $\hat{y}$ (and potentially other inputs such as ambient condition) and outputs a function $g(\hat{y})$. This network is pre-trained with labelled laboratory data. On the top right section of the figure (enclosed in black dashed line), a DeepONet model takes inputs of time $t$ and an input function $u$ to approximate a state $G(u)(t)$. This network is also pre-trained with labelled laboratory data. The physics/residual loss is shown in the bottom right part of the figure (enclosed within the blue dashed line). The PINN approximates the solution to differential equations. The total loss $\mathcal{L}(\bm{\theta})$ includes the loss of equation as well as the data. Weights $\lambda_{1}$ and $\lambda_{2}$ are applied to the data and physics loss respectively, and may be fixed or adaptive depending on the problem and solution method. $\bm{\theta}=\{\bm{W},\bm{b},\bm{\Lambda}\}$ represents the parameters of the PINN, where $\bm{W}$ and $\bm{b}$ are the weights and biases, and $\bm{\Lambda}$ are the unknown parameters of the ODE; $\sigma$ denotes the activation function used for the PINN model, and $r$ is the residual for the equation. Further, $\bm{\theta}^{P}=\{\bm{W}^{P},\bm{b}^{P}\}$ denotes the parameters of the pre-trained neural network $g(y)$. Similarly, $\bm{\theta}_{DP}=\{\bm{W}_{DP},\bm{b}_{DP}\}$ represents the parameters of the pre-trained DeepONet $G(u)(t)$.
@@
In section 1.2, we discussed the inverse problem of interest in this study. Our objectives are to: (1) learn the dynamics of the six gas flow dynamics state estimates, three of which are approximated using two DeepONets, and (2) infer the unknown parameters in the system, given field measurements of system variables $\{p_{im},p_{em},\omega_{t},W_{egr},T_{em}\}$ as well as Eqs. (3)-(10), using PINNs. To achieve these objective, we utilize four DNNs as surrogates for the solutions to different equations, as shown in Table 4. We encode Eqs (3)-(10) using the automatic differentiation and formulate the physics-informed loss. The total loss function considered for training the PINNs model is given by:
@@
$\begin{split}\mathcal{L}(\bm{\theta},\bm{\mathcal{\varLambda}},\bm{\lambda}_{p_{im}},\bm{\lambda}_{p_{em}},\bm{\lambda}_{\omega_{t}},\bm{\lambda}_{W_{egr}},\lambda_{T_{em}},\lambda_{T_{1}})=&\mathcal{L}_{p_{im}}+\mathcal{L}_{p_{em}}+\mathcal{L}_{\omega_{t}}+100\times\mathcal{L}_{x_{r}}+\lambda_{T_{1}}\times\mathcal{L}_{T_{1}}+\\ &\mathcal{L}^{ini}_{p_{im}}+\mathcal{L}^{ini}_{p_{em}}+\mathcal{L}^{ini}_{\omega_{t}}+100\times\mathcal{L}^{ini}_{x_{r}}+100\times\mathcal{L}^{ini}_{T_{1}}+\\ &\mathcal{L}^{data}_{p_{im}}(\bm{\lambda}_{p_{im}})+\mathcal{L}^{data}_{p_{em}}(\bm{\lambda}_{p_{em}})+\mathcal{L}^{data}_{\omega_{t}}(\bm{\lambda}_{\omega_{t}})+\\ &\mathcal{L}^{data}_{W_{egr}}(\bm{\lambda}_{W_{egr}})+\lambda_{Tm}\mathcal{L}^{data}_{T_{em}},\end{split}$ (20)
@@
Table 4: Neural network for PINN: Neural network surrogates considered to approximate the state variables and $x_{r}$ and $T_{1}$ for solving the inverse problems. $\mathcal{N}_{i}(t;\bm{\theta}_{i}),i=1,...,4$ denotes the surrogate for the $i^{th}$ DNN parameterized by $\bm{\theta}_{i}=\{\bm{W},\bm{b}\}_{i}$. The input to these DNN is time $t$. DNN $\mathcal{N}_{1}(t;\bm{\theta}_{1})$ has two outputs for $\{p_{im}$ and $p_{em}\}$; the remaining DNNs have only one output.
@@
Output scaling ×$\times 10^{5}$$\times 1$$\times 300$$\times 5.0\times 10^{3}$

### v2

@@ 2.1 Overview
    Empirical relationships: We consider various empirical relationships between variables in the engine model. In prior study, [30] demonstrated how these empirical formulae can be approximated using deep neural networks (DNN). The empirical formulae are engine-specific, and DNNs can provide a generalized representation when the exact relationships are unknown by training the DNN with data from the specific engine. In addition, to account for the systematic (e.g. manufacturing) variations of the engine, we introduce stochastic variations into the empirical relationship by using dropouts in DNN.
@@
    Independent States Approximation: We note that the states describing the EGR and VGT dynamics are independent of the rest of the states, with no unknown parameters in these dynamics. We approximated these states using two DeepONets, one for the two EGR states ($\tilde{u}_{egr1}$ and $\tilde{u}_{egr2}$) and one for the VGT state ($\tilde{u}_{vgt}$) as discussed in section 2.3. These DeepONets are pre-trained using simulation data and used as surrogates to predict the actuator states.
@@
    Hybrid method: We propose a hybrid method combining the pre-trained DNN, the pre-trained DeepONet and PINN model, as detailed in section 2.4. We also use self-adaptive weights [36] for improved convergence in all examples of the proposed methods.
@@
    Transfer Learning Strategies: Lastly, to reduce the computation cost (in terms of the number of epochs and trainable parameters), we consider two TL strategies:
@@
      a hybrid PINN multi-stage transfer learning method, discussed in section 2.5,
@@
Furthermore, as noted above, we use pre-trained networks for empirical formulae (DNNs) and for EGR and VGT states (DeepONets). These networks are trained using simulated data (we label this as laboratory data for nomenclature simplicity), which corresponds to the design value of the unknown parameters. However, in field, the values of the unknown parameters may change due to natural systematic changes over time. Thus, we test our proposed hybrid PINN method on a set of parameters different from the design parameters used for generating laboratory data for training our empirical networks and DeepONets. This helps us validate whether our hybrid PINN model can accurately identify parameters independently of the predictions from the pre-trained networks, thereby allowing us to analyze any potential bias originating from this hybrid approach. The different values of the unknown parameters considered for generating laboratory and field data are shown in Table 1.
Table 1: True value of estimated parameters considered for generating laboratory and field data conditions. The “Laboratory data” is considered for training the DNNs for empirical formulae, DeepONets for EGR and VGT actuator dynamics, and training the body for few-shot TL. “Field data” represents data used for training the hybrid PINN framework for parameter estimation, mutli-stage TL and few-shot TL. For confidentiality, we represent the true values for field parameter values as 1, and the laboratory data are scaled with respect to field data. The “Mask function” ensures the unknown parameters are always positive, and “Scaling” factor enables the conversion of parameters back to the physical domain.
$A_{egrmax}$ $A_{vgtmax}$ $\eta_{sc}$ $h_{tot}$
Laboratory data $1.333$ $1.1663$ $0.9179$ 11
Field data 11 11 11 11
Mask and scaling considered for PINN model
Mask function $\exp(.)$ $\exp(.)$ softplus(.) $\exp(.)$
Scaling $\times 10^{-4}$ $\times 10^{-4}$ $\times 1$ $\times 10$
@@ 2.2 Neural network surrogates for empirical formulae
The empirical equations are discussed in section A
@@
where the subscript $i=1,2,\dots,6$ indicate different neural networks shown in Table 2, $n_{i}$ represents the number of labelled training data for $i$th neural network. $\bm{x}_{i}$ denotes the input corresponding to the $i$th network, $\hat{y}_{i}$ and $y_{i}$ represent the estimated output and the corresponding labelled values, respectively for the $i$th network. The laboratory data used for training each neural network have been discussed in previous work [30] along with detailed calculations for each labelled data. We also consider a 3% Gaussian noise in the simulated data of the output variables during training to account for the field measurement conditions. These networks are trained using the Adam optimizer. After training, we integrate these DNNs into our hybrid PINN framework to replace the empirical models. These pre-trained neural networks are shown with the output g(y^)g($\hat${y}) in Fig. 2 and Fig. 4 in sections 2.4 and 2.6.
We apply a 1% dropout for the hidden layers of these networks. Notably, dropout is employed not only during the training of these networks but also during the PINN optimization process for the inverse problem and prediction. This helps us to analyze the potential variation in PINN estimates expected due to systematic (e.g., manufacturing) variability for real systems. We also added a 3% Gaussian noise to the output state variables during training to replicate the expected noise in laboratory measurements.
@@ 2.3 DeepONets for EGR and VGT states
Based on the universal approximation theorem of operators [37], Lu et al. [32] proposed the Deep Operator Network (DeepONet), which maps an infinite-dimensional function ($u\in\bm{U}$) to another infinite dimensional output function ($v\in\bm{V}$) in the Banach space. The conventional DeepONet consists of two DNNs, known as branch and trunk networks. The trunk network takes the spatio-temporal information (e.g. time $t$, domain points x,y,z{x,y,z}) as input and provides the bases function as output. The branch network takes the input function ($u\in\bm{U}$) at fixed points and approximates the coefficients for each basis function. Importantly, the DeepONet architecture does not impose restrictions on the type of networks used for trunk and branch networks. The ability of DeepONet to learn complex mapping between input and output functional spaces makes it suitable for various applications in scientific machine learning. Mathematically, DeepONet learns an operator that maps $\mathcal{G}_{\bm{\theta}}:\bm{U}\rightarrow\bm{V}$, and the output is given as
@@
We note here that the engine is a continuous dynamical system, and this makes it challenging to identify proper operating sub-spaces for input-output mapping for the operator network learning. Hence, it is important to study the driving behaviour (engine running condition) over large duration and sample sizes. An appropriate window size (time duration) is required in which correct input-output function spaces can be assumed while preserving the dynamic behaviour of the system. Here, we consider a 60-sec window and approximate the independent state variables of the engine ($\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$, $\tilde{u}_{vgt}$) using Causality-DeepONet [39], as proposed by Liu et al. [39]. Causality-DeepONet integrates causality and convolution of the input function in its formulation, which enhances accuracy in approximating oscillatory dynamic problems with a smaller training sample set. We find that this framework is particularly useful in capturing the rapidly changing dynamics of a diesel engine signal.
In this study, we use Causality-DeepONets to approximate the dynamics of the EGR actuator dynamics, i.e., $\tilde{u}_{egr1}$ (Eq. (6)) and $\tilde{u}_{egr2}$ (Eq. (7)) and the VGT actuator dynamics $\tilde{u}_{vgt}$ (Eq. (8)). Details of the two DeepONets are shown in Table 3. From hereon, we refer to Causality-DeepONet simply as DeepONet unless stated otherwise. These DeepONets approximate solutions to non-homogeneous initial value problems by learning the operator mapping between input function ($u$) and the initial condition ($\tilde{u}_{0}$) to the output ($\tilde{u}(t))$ using individual branch networks for the input function and initial condition, as described in Table 3. We note that the equations governing EGR and VGT actuator dynamics are first-order differential equations that can be computed efficiently using fast numerical methods. However, discrepancies between the mathematical model and experimental data were observed in [31]. By employing DeepONets trained on historical data, we can obtain a better representation of real system dynamics as compared to a numerical method that relies on physical equations. Hence, our approach of using DeepONet surrogates for predicting actuator dynamics is likely to show better correlation when used on real system. We can write the operator mapping from input function ($u$) and initial condition ($\tilde{u}_{0}$) to the dynamics of the output ($\tilde{u}(t)$) as
We have discussed the DeepONet architecture, training, and numerical results in Appendix E
@@
DeepONet $\mathcal{D}_{1}(t,\bm{x};\bm{\theta}_{DP1})$ $\mathcal{D}_{2}(t,\bm{x};\bm{\theta}_{DP2})$
No. of branch 4 2
Branch Input ($\bm{x})$ $\{u_{degr}$, $\tilde{u}_{egr1,0}\}$, $\{u_{degr}$, $\tilde{u}_{egr2,0}\}$ $u_{dvgt}$, $\tilde{u}_{vgt,0}$
Output Variable $\{\tilde{u}_{egr1}$, $\tilde{u}_{egr2}\}$ $\{\tilde{u}_{vgt}\}$
Eq. (A.16) and (A.17) (A.25)
For equations, refer Appendix.
For details on DeepONet architecture, training, and results, refer Appendix E
@@ 2.4 A hybrid model using DeepONet and PINN for the engine problem
In the previous section (2.3), we discussed how the three states corresponding to EGR and VGT can be approximated using DeepONet surrogates. In section 2.2, we introduced how we use DNNs to estimate empirical relationships in the engine model. In this section, we discuss in detail how we construct a hybrid model combining DeepONet, empirical DNNs, and PINN to solve the inverse problem of model-driven parameter identification. We propose a hybrid PINN model combining DeepONet and DNN to predict the gas flow dynamics and estimate unknown parameters. Preliminary background on use of PINNs to approximate the state variables of an ODE is presented in Appendix B
@@
where $\bm{\theta}$ denotes the set of parameters of the neural network, i.e. the weights, $\bm{W}$ and the biases, $\bm{b}$; $\bm{\mathcal{\varLambda}}$ are the unknown scalar parameters to be estimated. The parameters $\lambda_{1}$ and $\lambda_{2}$ are the weights which control the convergence of each loss component. These $\lambda$ values may be constant or self-adaptive [36] depending on the problem setup and solution method. The data loss may be considered as the mean squared error (MSE) loss of the known labelled data, given by:
@@
where $N$ is the number of collocation points chosen for solving the equation and $r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})$ represents the residual of the corresponding differential equation. The residual $r(t_{i};\bm{\theta};\bm{\mathcal{\varLambda}})$ should ideally be zero throughout the entire domain (or time duration). The objective, therefore, is to obtain the optimal parameters ($\bm{\theta}$) of the neural network as well as the unknown parameters ($\bm{\mathcal{\varLambda}}$) in the system. We minimize the loss function Eq. (14) to obtain the optimal values of these parameters. Furthermore, we also incorporate self-adaptive weights [36] in the loss. We minimize the loss function with respect to the parameters ($\bm{\theta}$ and $\bm{\varLambda}$) and maximize with respect to the self-adaptive weights ($\lambda$). Thus, the optimization process may be written as,
@@
To integrate the complete hybrid model, we combine the DNN models used for approximating the empirical formulae as discussed in section 2.2 and the two DeepONet models for approximating actuator states $\tilde{u}_{egr1}$, $\tilde{u}_{egr2}$ and $\tilde{u}_{vgt}$ with the PINN DNN model, as shown in Fig. 2. We fix the parameters of DeepONets and DNNs (for the empirical formulae) when training the hybrid PINN model. Eq. (14), demonstrates how PINNs can be used for solving one differential equation. For systems involving multiple equations, one can either use one DNN with multiple outputs or multiple DNNs each with one output. The loss function in Eq. (14) can be extended to include the loss function for all the equations and data by summing up individual loss functions for each equation and data. This hybrid approach leverages pre-trained networks and a PINN network that can be trained on live data for efficient and accurate estimation of engine parameters and state variables.
@@
Fig. 2: Schematic of a hybrid model of DeepONet and physics-informed neural networks (PINNs) for inverse problems. In the left part of the figure, a PINN approximates the solution $y$ to a differential equation with time $t$ as input. The top left section (enclosed within black hatched lines) represents another DNN which takes as input $\hat{y}$ (and potentially other inputs such as ambient condition) and outputs a function g(y^)g($\hat${y}). This network is pre-trained with labelled laboratory data. On the top right section of the figure (enclosed in black dashed line), a DeepONet model takes inputs of time $t$ and an input function $u$ to approximate a state $G(u)(t)$. This network is also pre-trained with labelled laboratory data. The physics/residual loss is shown in the bottom right part of the figure (enclosed within the blue dashed line). The PINN approximates the solution to differential equations. The total loss $\mathcal{L}(\bm{\theta})$ includes the loss of equation as well as the data. Weights $\lambda_{1}$ and $\lambda_{2}$ are applied to the data and physics loss respectively, and may be fixed or adaptive depending on the problem and solution method. $\bm{\theta}=\{\bm{W},\bm{b},\bm{\Lambda}\}$ represents the parameters of the PINN, where $\bm{W}$ and $\bm{b}$ are the weights and biases, and $\bm{\Lambda}$ are the unknown parameters of the ODE; $\sigma$ denotes the activation function used for the PINN model, and $r$ is the residual for the equation. Further, $\bm{\theta}^{P}=\{\bm{W}^{P},\bm{b}^{P}\}$ denotes the parameters of the pre-trained neural network $g(y)$. Similarly, $\bm{\theta}_{DP}=\{\bm{W}_{DP},\bm{b}_{DP}\}$ represents the parameters of the pre-trained DeepONet $G(u)(t)$.
@@
In section 1.2, we discussed the inverse problem of interest for this study. Our objectives are to: (1) learn the dynamics of the six gas flow dynamics state estimates, three of which are approximated using two DeepONets, and (2) infer the unknown parameters in the system, given field measurements of system variables $\{p_{im},p_{em},\omega_{t},W_{egr},T_{em}\}$ as well as Eqs. (3)-(10), using PINNs. To achieve these objectives, we utilize four DNNs as surrogates for the solutions to different equations, as shown in Table 4. We encode the physics loss corresponding to Eqs (3)-(10) (EGR, VGT actuator equations are not included as these are approximated using DeepONets) using automatic differentiation and define the physics-informed loss as:
@@
$\begin{split}\mathcal{L}(&\bm{\theta},\bm{\mathcal{\varLambda}},\bm{\lambda}_{p_{im}},\bm{\lambda}_{p_{em}},\bm{\lambda}_{\omega_{t}},\bm{\lambda}_{W_{egr}},\lambda_{T_{em}},\lambda_{T_{1}})\\ =&\mathcal{L}_{p_{im}}+\mathcal{L}_{p_{em}}+\mathcal{L}_{\omega_{t}}+100\times\mathcal{L}_{x_{r}}+\lambda_{T_{1}}\times\mathcal{L}_{T_{1}}+\\ &\mathcal{L}^{ini}_{p_{im}}+\mathcal{L}^{ini}_{p_{em}}+\mathcal{L}^{ini}_{\omega_{t}}+100\times\mathcal{L}^{ini}_{x_{r}}+100\times\mathcal{L}^{ini}_{T_{1}}+\\ &\mathcal{L}^{data}_{p_{im}}(\bm{\lambda}_{p_{im}})+\mathcal{L}^{data}_{p_{em}}(\bm{\lambda}_{p_{em}})+\mathcal{L}^{data}_{\omega_{t}}(\bm{\lambda}_{\omega_{t}})+\\ &\mathcal{L}^{data}_{W_{egr}}(\bm{\lambda}_{W_{egr}})+\lambda_{Tm}\mathcal{L}^{data}_{T_{em}},\end{split}$ (18)
[…]
@@
Table 4: Neural network for PINN: Neural network surrogates considered to approximate the state variables and $x_{r}$ and $T_{1}$ for solving the inverse problems. $\mathcal{N}_{i}(t;\bm{\theta}_{i}),i=1,...,4$ denotes the surrogate for the $i^{th}$ DNN parameterized by $\bm{\theta}_{i}=\{\bm{W},\bm{b}\}_{i}$. The input to these DNN is time $t$. DNN $\mathcal{N}_{1}(t;\bm{\theta}_{1})$ has two outputs for $\{p_{im}$ and $p_{em}\}$; the remaining DNNs have only one output. The multihead networks are pre-trained offline. During multi-stage transfer learning the number of parameters changes with epoch as discussed in the
@@
Hidden layer activation Tanh Tanh Tanh Tanh
@@
Output scaling $\times 10^{5}$ $\times 1$ $\times 300$ $\times 5.0\times 10^{3}$
@@
Network size
Multi-stage TL [1, 15, 15, 15, 2] [1, 15, 15, 15, 1] [1, 15, 15, 15, 1] [1, 15, 15, 1]
Multihead network $[1,250,250,250,125,480]$ $[1,170,170,170,50,240]$ $[1,50,50,20,240]$ $[1,120,120,70,240]$
During few-shot TL $[125,1]$ and $[125,1]$ $[50,1$] $[20,1]$ $[70,1]$